### v1

# Geospatial analysis of toponyms in geotagged social media posts

Abstract
Place names, or toponyms, play an integral role in human representation and communication of geographic space. In particular, how people relate each toponym with particular locations in geographic space should be indicative of their spatial perception. Here, we make use of an extensive dataset of georeferenced social media posts, retreived from Twitter, to perform a statistical analysis of the geographic distribution of toponyms and uncover the relationship between toponyms and geographic space. We show that the occurrence of toponyms is characterized by spatial inhomogeneity, giving rise to patterns that are distinct from the distribution of common nouns. Using simple models, we quantify the spatial specificity of toponym distributions and identify their core-periphery structures. In particular, we find that toponyms are used with a probability that decays as a power law with distance from the geographic center of their occurrence. Our findings highlight the potential of social media data to explore linguistic patterns in geographic space, paving the way for comprehensive analyses of human spatial representations.
I Introduction
When we speak or write about geographic spaces, we generally communicate them using the names of places, or toponyms. Although any point on Earth can be specified by geographic coordinates, one would rarely refer to a place in day-to-day conversation as, for example, “35.7°35.7°35.7  °N, 139.7°139.7°139.7  °E”; instead, we would usually represent it by a toponym, such as “Tokyo”. The use of toponyms reflects the way we perceive and mentally structure geographic space. Unlike geographic coordinates, which are objective and unambiguous, the area a toponym refers to is often vague and difficult to define; even for the names of administratively defined areas (such as municipalities), how people use them in colloquial settings may have only a loose correspondence with the officially demarcated boundaries [1, 2]. This however does not mean that the extent that each toponym denotes can be arbitrarily defined by individuals; since the main function of toponyms is to effectively communicate geographic information, there must be a shared consensus within the population that determines the areas they represent.
The objective of this study is to understand such a spontaneous and collective relationship between geographic space and the use of toponyms. In order to quantitatively examine this relationship, one needs to collect large scale data on the locations people associate with each toponym. The abundance of user-generated content online, especially in social media, offers a promising opportunity for this purpose. In particular, Twitter (rebranded to “X” in 2023), one of the major social media platforms, had provided free access to posts on the platform through APIs until 2023, allowing researchers to perform statistical analysis and find patterns and trends in the data. On Twitter, users could opt to attach to each of their post geolocation metadata (a geotag) that represent the geographic coordinates of the GPS location of their device. These user-annotated geotags as well as the content of the posts allow us to explore the spatial dimension of user behavior and language use. For instance, by leveraging the fact that the set of vocabulary that appears in the text of posts varies according to the whereabouts of users, studies have found that toponyms in text can be disambiguated [3] and the location of individual users can be identified [4, 5]. When aggregated at the population level, the data can be used to study dialectical variations and are of interest to sociolinguistics and linguistic geography [6, 7, 8, 9, 10].
We note that the geotag attached to a post does not necessarily correspond to the place referred to in the text [11, 12]. This discrepancy can arise from inaccurate tagging [13], but it can also be a manifestation of the geographic awareness and identity of the user, i.e., how users perceive and choose to represent their location or the place under discussion [14, 15, 16]. Users may (mis)represent their location for various reasons, such as privacy concerns, a desire to be associated with a particular place, or simply because they are referring to a location that is not their own.
In this study, we focus on understanding the geographic distribution of geotagged posts and the toponyms they contain. Our goal is not to identify patterns of toponym usage at the individual user level, but rather to understand the collective, population-level knowledge about the relationship between toponyms and geographic locations. Similar questions have been addressed in the literature: Hollenstein and Purves [17] used user-annotated geotag data from the image hosting service Flickr to delineate city centers and neighborhoods; Hu and Janowicz [18] studied the names of points of interest (such as restaurants and shops) registered on Yelp, a user-generated local business review platform, in metropolitan areas in the United States. These studies were primarily focused on urban areas, and therefore the analysis was limited to a relatively small geographic scale. In addition, the results were rather descriptive and were not aimed at deriving general laws of toponymic occurrence. In contrast, we use a data-driven modeling approach to uncover the underlying principles governing the occurrence of toponyms of different granularities on a larger geographic scale, aiming to understand how these distributions reflect collective spatial perceptions.
We first show that the distribution of the total set of geotagged posts is characterized by a power-law density distribution, reflecting the inhomogeneity of the population distribution. We then focus on the geographic distribution of individual toponyms. We find that the occurrence of each toponym is inhomogeneously distributed in geographic space, in contrast to the distributions of common nouns and toponyms referring to a place outside the observation area. We show that such inhomogeneity cannot be explained without taking into account the spatial specificity of each toponym, and propose a simple alternative model that reproduces features of the observed data. We conclude by discussing the implications of our work.
II Material
[…]
As of 2019, Japan is divided into 47 prefectures as the first level of administrative division, and $1741$ municipalities (cities, towns, villages, and twenty-three special wards of Tokyo) as the second level of division. In addition, some large cities, referred to as “designated cities”, have administrative, non-autonomous subdivisions known as wards. Prefectures are sometimes grouped into seven to thirteen regions, although regions are not official administrative units, and the name and extent of each region can be ambiguous.
From the full sample of the $277$ million geotagged posts, we extract the subset of posts that contains each of 24 Japanese toponyms that refers to regions, prefectures, cities, and wards (special wards of Tokyo Metropolis and wards of designated cities) in Japan. The list of toponyms sampled in this work and the locations of administrative divisions they refer to can be found in Fig 1. In addition, we extracted posts that contain twelve common Japanese nouns and six Japanese toponyms that refer to places outside Japan. We call these samples for individual keywords (toponyms and nouns) as keyword subsamples. In the following, we use only the information about the number of posts tagged inside each grid cell for each sample, and disregard the content of the text or user information. Refer to Table 1 for the notation and definition of variables used in this work.
Figure 1: Places in Japan denoted by the toponyms studied in this paper. In this study, we sample 24 toponyms: the names of (A) six regions, (B) six prefectures, (C) six major cities, and (D) six wards (submetropolitan/submunicipal districts). The colored areas in each panel show the administratively defined geographic area (except for cities) denoted by each toponym. (A) The extent of each region is not uniquely defined. Here we show one of the commonly used classification of regions. (C) The colored area shows the metropolitan employment area [19], which is considered to be more representative of urban activity than administratively defined city areas. Note that Kyoto, Hiroshima, and Fukuoka are used both as the names of the cities and as the names of the prefectures of which the cities are the capitals. (D) Prefectural boundaries are shown for visual guidance.
Table 1: Variables used in this work. Subscript $c$ for the grid cell index may be omitted when it is clear.
Symbol Definition
$w$ Keyword on which the dataset is subsampled
$c$ Index of grid cell
$n_{\mathrm{all},c}$ Number of all posts tagged to grid cell $c$
$n_{w,c}$ Number of posts containing $w$ and tagged to grid cell $c$
$N_{w}$ Total number of geotagged posts containing keyword $w$; $N_{w}=\sum_{c}n_{w,c}$
$A_{c}$ Area of grid cell $c$
$\sigma_{\mathrm{all},c}$ Density of all posts tagged to grid cell $c$; $\sigma_{\mathrm{all},c}:=n_{\mathrm{all},c}/A_{c}$
$\sigma_{\mathrm{res},c}$ Resident population density in grid cell $c$
$\sigma_{\mathrm{emp},c}$ Density of employed population (number of personnel working at business sites) in grid cell $c$
$\sigma_{w,c}$ Density of posts containing keyword $w$ and tagged to grid $c$; $\sigma_{w,c}:=n_{w,c}/A_{c}$
$\phi_{w,c}$ Occurrence ratio of keyword $w$ among all posts tagged to grid cell $c$; $\phi_{w,c}:=n_{w,c}/n_{\mathrm{all},c}$
$p_{w,c}$ (Latent) probability that keyword $w$ is contained in a post tagged to grid cell $c$
III Results
III.1 Spatial distribution of geotagged posts
Let us first study the geospatial distribution of the full sample of geotagged posts before looking at the distribution of each toponym subsample (Fig 2). The geolocations to which the sampled posts are tagged are not uniformly distributed within the observed area, as shown in Fig 2A. A large number of posts are concentrated in relatively few grid cells in large metropolitan areas, such as the city centers of Tokyo and Osaka, while few posts are found in most of the grid cells. The heterogeneity of the spatial distribution of geotagged posts is also evident from the heavy-tailed probability distribution of the number of posts sampled in each grid cell (Fig 2D). Namely, it is characterized by two power laws with different exponents: the bulk part is characterized by an exponent of approximately $-1.2$ while the tail part follows a steeper power law with an exponent of about $-2.8$.
Figure 2: The spatial densities of geotagged posts, resident population, and employed population. (A–C) Geographic distribution of the three densities. Note that the population data are geographically limited inside Japan, while the geotagged posts are sampled in bounding box of Japan, which also includes neighboring countries. (D) Probability distribution of density (the number of geotagged posts per unit area). (E, F) Scatter plots showing the correlation between each of the population densities and the geotagged post density. Pearson and the Spearman correlation coefficients are shown below the plot. (G) Probability distributions of the three densities, each rescaled by its mean.
To investigate the origin of this heterogeneous distribution, we compare the geotagged post statistics with census data for the resident population and the employed population. The employed population is defined as the number of permanent or temporary employees, self-employed individuals, contractors, and unpaid staff in family-owned businesses, whose main working site is located in each grid cell. From Figs. 2B and  2C, one can see the similarity in the spatial distribution between the densities of geotagged posts, resident population, and employed population. In fact, the geotagged post density is strongly correlated with the population densities (Figs. 2E,  2F). The correlation with the post density is stronger for the employee population density than for the resident population density, both in terms of Pearson’s correlation and Spearman’s rank correlation. This may suggest that the social media posts are made more in the daytime than at night. Figure 2G shows the probability density functions of the geotagged post density, the resident population density, and the employee population density collapse to one another when they are scaled by the average density, strongly suggesting that the uneven distribution of population gives rise to the heterogeneity in the spatial distribution of geotagged posts.
III.2 Spatial distributions of toponym subsamples
We now turn to our main question: How are geotagged posts containing each toponym, i.e., each toponym subsample, spatially distributed? To address this question, we show the results for “Fukuoka”, a toponym referring to the sixth largest city in Japan as of 2019, located in the southwestern part of the country, as an illustrative example.
Figure 3A shows the spatial distribution of the occurrence of “Fukuoka”, while Fig 3C shows the probability density functions of the gridwise occurrence density $\sigma_{w}$. Both figures suggest that the “Fukuoka” occurs heterogeneously across different grid cells. The same observation can be made for other toponyms, as shown in Fig S2 in Supporting Information. It is noteworthy that toponym subsamples of different sizes follow the same scaling, which implies that the spatial heterogeneity does not depend on the popularity of toponyms. On the other hand, each distribution is characterized by a single power law, which is a clear deviation from the scaling observed for the full sample.
Figure 3: Occurrence pattern of “Fukuoka”. (A) Geographic distribution of density $\sigma_{w}$. (B) Geographic distribution of occurrence ratio $\phi_{w}$. (C) Probability distributions of spatial density for all geotagged posts (in gray) and for “Fukuoka” (in blue). (D) Scatter plot showing the relationship between the total number of geotagged posts $n_{\mathrm{all}}$ and the number of posts containing “Fukuoka” $n_{w}$ in each grid cell. The lines represent contours along which the density of points (kernel density estimate) on the double logarithmic scale is constant. The region inside each contour, from dark to light colors, contains 90.0%, 99.0%, and 99.9% of the data points, respectively. (E) The same scatter plot as panel D, but with points colored by the distance between the grid cell and the center $O_{w}$.
While the spatial density (occurrence per unit area) of posts containing the word “Fukuoka” is high in Fukuoka and neighboring areas, it is also high in other large metropolitan areas such as Tokyo and Osaka, which are geographically distant from Fukuoka. However, this is presumably just by chance due to the large total number of posts tagged in these areas. To account for the difference in total sample size $n_{\mathrm{all}}$ in each grid cell, we normalize the toponym subsample size $n_{w}$ by $n_{\mathrm{all}}$ and obtain the fraction of occurrence of the toponym. We clearly see that posts contain word “Fukuoka” with higher probabilities in the region around the city of Fukuoka (Fig 3B).
To further understand the relationship between the full sample and the toponym subsamples, we create a scatter plot for each toponym as in Fig 3D, where the horizontal axis represents $n_{\mathrm{all}}$ and the vertical axis corresponds to $n_{w}$. Apart from the general trend that $n_{w}$ increases with $n_{\mathrm{all}}$, we see that there are two distinct branches of scaling, with one increasing faster than the other. That is, $n_{w}$ increases as a function of $n_{\mathrm{all}}$ in two different ways. This two-branch scaling behavior is not specific to “Fukuoka” but is widely seen for different toponyms across various degrees of popularity and granularity (Fig S3 in Supporting Information). In contrast, keyword subsamples for common nouns such as “wallet” and toponyms that refers to places outside the observed area such as “Hawaii” do not exhibit heterogeneity in spatial distribution, and the relationship between $n_{\mathrm{all}}$ and $n_{w}$ only shows a single scaling (see Figs. 4B and 4D; see also Fig S3 for the scatter plots for all keywords sampled in this study). This suggests that the two-branch scaling is unique to the toponyms that refer to places within the observed area and is not present in other types of words. Naturally, we expect that such pattern stems from the geospatial specificity of toponym use.
In particular, we hypothesize that one of the two scaling branches observed for toponym subsamples corresponds to the distribution of the toponym that is spatially specific to the area it refers to, while the other branch represents the use of the toponym that is not spatially specific, similar to that of common nouns. To test this hypothesis, we first need to identify the area that the toponym refers to. Instead of relying on external sources such as gazetteers, we do this in a data-driven way: we define the center of the geographic distribution of toponym $w$ as the grid cell with the highest frequency of $w$:
$O_{w}=\operatorname*{argmax}_{c\in S_{w}}\,n_{w,c},$ (1)
where $S_{w}=\{c\mid\phi_{w,c}\geq\epsilon\}$ is the set of cells with an occurrence ratio equal to or greater than $\epsilon$. This condition prevents a cell from being selected as the center merely due to the large total number of posts. Here we set $\epsilon=0.01$, i.e., toponym $w$ must appear in at least 1% of all posts tagged to a cell for the cell to be included in $S_{w}$. As we can see in Fig 3E, the grid cells that constitute the two branches are clearly distinct in terms of distance from the center of the toponym subsample. The branch with faster scaling is geographically closer to the center while the branch with slower scaling is relatively far from the center. This corroborates our hypothesis that the two-branch scaling arise from the spatial specificity of the toponym.
III.3 Location-independent model
The results in the previous section give us an intuitive understanding of the geographic distribution of toponyms. In the next two sections, we present a model-based analysis to validate our intuition and to characterize the empirical observation in a quantitative way. Specifically, we introduce a model where whether each geotagged post contains word $w$ is a Bernoulli trial with success probability $p_{w,c}$, which possibly depends on the location $c$ it is tagged to. The number $n_{w,c}$ of posts in grid cell $c$ that contains word $w$ out of $n_{\mathrm{all},c}$ total posts follows a binomial distribution:
$P(n_{w,c}\mid n_{\mathrm{all},c},\,p_{w,c})=\binom{n_{\mathrm{all},c}}{n_{w,c}}p_{w,c}^{n_{w,c}}(1-p_{w,c})^{n_{\mathrm{all},c}-n_{w,c}}.$ (2)
This model essentially posits that the underlying mechanism of the occurrence of a word can be summarized by probability $p_{w,c}$ specific to each grid cell $c$. Importantly, it is based on the simplifying assumption that the correlations between occurrences of different words can be neglected.
We first examine if the empirical pattern can be explained by the simplest version of this binomial model which assumes that $p_{w,c}=p_{w}$ for all grid cell $c$; that is, the toponym occurs with a constant probability independent of location. The unbiased and maximum likelihood estimator for $p_{w}$ can be obtained simply by dividing the size of the toponym subsample by the number of all geotagged posts: $\hat{p}_{w}=N_{w}/N_{\mathrm{all}}$.
Figure 4A shows the distribution of occurrence of “Fukuoka” against $n_{\mathrm{all}}$ in empirical data and the expectation from the location-independent binomial model. We observe that the empirical distribution is not consistent with the model. In particular, the model does not reproduce the two-branch scaling behavior seen in the empirical data and exhibits a single scaling instead. We confirm the same kind of discrepancy for all the domestic toponyms we studied (see Fig S4 in Supporting Information for the full results).
Figure 4: Relationship between $n_{w}$ and $n_{\mathrm{all}}$ in empirical and model distributions. (A) Scatter plots for different keywords. (B, C) Kernel density profiles of empirical and model distributions. In each panel, the red contours show the density profile obtained from the location-independent model (B) or the core-periphery model (C), fitted to the empirical occurrence pattern of each word, represented by the blue contours. As in Fig. 3D, the region inside each contour, from dark to light colors, contains 90.0%, 99.0%, and 99.9% of the data points, respectively. Note that the scatter plot of empirical data and its density profile for “Fukuoka” are identical to those in Fig 3D.
The implication of this observation becomes apparent through juxtapositions against patterns for common nouns and foreign nouns. The location-independent model shows a close agreement with the empirical data for “wallet” (Fig 4B), implying that the word indeed occurs at a constant rate in every grid cell. In general, we find that the empirical data and the location-independent model are in fairly good agreement for many common nouns that refer to objects (e.g., “telephone”) or abstract concepts (e.g., “society”). For a comparison between the model and the empirical distributions for these words, see Fig S4 in Supporting Information. On the other hand, there is a large deviation between the empirical and model distributions for the word “airport” (Fig 4C), suggesting that the occurrence at a constant rate is not a shared characteristic among all common nouns. This can be explained by the fact that common words such as “airport” are often used in combination with toponyms (e.g. “Narita Airport”), and are therefore more likely to be used in specific places. Finally, for foreign toponyms, such as “Hawaii” shown in Fig 4D, we observe the single scaling pattern in both the empirical and model distributions, although the empirical distributions generally show slightly greater variance than the model. While these toponyms are not semantically associated with a specific place in the observed area, they may occur with higher probability around international airports from which people travel to the places these toponyms refer.
Beyond visual inspection, the discrepancy between the location-independent model $P$ and the data can be quantified using relative entropy, also known as the Kullback–Leibler (KL) divergence. Intuitively, relative entropy measures the dissimilarity from one probability distribution to another. Here, we wish to quantify the dissimilarity of the empirical data from the model, each of which is a probability distribution on the total number of posts in each cell and the number of posts with word $w$ in each cell. Specifically, we employ a modified version of relative entropy, denoted by $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$, that takes into account the cells with at least one occurrence of $w$. We elaborate on this particular definition and provide the rational for our choice in Supporting Information S1.
As shown in Fig 5, relative entropy $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$ clearly differentiates domestic toponyms and place nouns from foreign toponyms and common nouns without place connotations. All domestic toponyms except “Kanto” and all place nouns (“park”, “university”, “hotel”, “airport”, and “shrine”) are characterized by relatively large values of relative entropy, namely $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})>4$, implying that the empirical distribution is highly dissimilar from the location-independent binomial model for these words. Common nouns such as “trip” and “vegetable” are on the borderline ($D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})\simeq 4$); relative entropy for other common nouns and foreign toponyms are significantly smaller.
Figure 5: Dissimilarity of empirical data from the location-independent model. The dissimilarity is evaluated by relative entropy $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$. For each word, 3000 grid cells are randomly sampled 50 times. The error bar shows the 95% confidence interval.
The comparison with the location-independent model reveals how sensitive and specific the occurrences of toponyms and common nouns are to geographic locations. For toponyms, the spatial specificity may be readily observable by visualizing the geographic distribution of occurrence ratio $\phi_{w}$, as in Fig 3B. However, identifying location-specific common nouns may be more subtle, as these nouns are usually associated not with a single place but with multiple places across the observed area. As a result, their geographic distributions may appear visually indistinguishable from those of non-place nouns. In such cases, quantifying the dissimilarity from the binomial model through $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$ can serve as a good indicator of the geospatial specificity of word occurrence.
III.4 Core-periphery model
So far we have shown that the location-independent binomial model cannot account for the large variance and two-branch scaling behavior seen in empirical distribution for domestic toponyms. Taking a step towards reality while maintaining simplicity, we consider a location-dependent variant of the binomial model. Namely, we assume that the value of $p_{w,c}$ in Eq (2) can vary for different grid cell $c$.
If the probability $p_{w,c}$ that a post contains toponym $w$ is to be estimated independently for each grid cell $c$, its unbiased, maximum likelihood estimator would be equal to the occurrence ratio $\phi_{w,c}=n_{w,c}/n_{\mathrm{all},c}$. However, since this model has as many parameters as the number of grid cells, it provides little insight into the spatial patterns of toponym occurrence. Here we propose a simpler model that assumes the presence of a core, where the toponym occurs most frequently, and periphery, geographically distant grid cells with lower $p_{w,c}$. We naturally expect that the core indicates the geographical extent to which the toponym refers.
Plotting the occurrence ratio $\phi_{w,c}$ against the geodesic distance $d_{w,c}$ between the center $O_{w}$ and cell $c$, we observe an overall decreasing trend with distance from the center (Fig 6A). In fact, the binned averages of $\phi_{w,c}$ as a function of distance from the center imply a power-law decay of the average occurrence ratio at locations far from the center. Moreover, by grouping toponyms according to administrative level, as shown in Fig 6B, one can see that the onsets of the power law differ according to the granularity of each toponym. For toponyms referring to higher-level administrative units, which are typically larger in area, the power-law decay starts at a larger distance, while the average occurrence ratios are relatively stable at smaller distances. Conversely, toponyms denoting small administrative units, such as wards, exhibit power-law behavior starting at short distances with no noticeable plateau regime.
Figure 6: Core-periphery patterns of toponym occurrence. (A) Occurrence ratio $\phi_{w}$ of “Fukuoka” against distance $d_{w}$ from center $O_{w}$ (small black dots), overlaid with the average for each logarithmic bin (red circles). The solid and dashed lines represent the maximum likelihood fits of the location-independent and core-periphery binomial models. (B) Average occurrence ratio as a function of $d_{w}$ for all the domestic toponyms studied in this work. (C) Maximum likelihood estimator of the core-periphery model parameters for each toponym. We represent estimated radius $\hat{r}_{w}$ by bars colored according to the category of the toponym (lower axis) and estimated exponent $\hat{a}_{w}$ by gray circles (upper axis). (D) The fitted core-periphery model compared to the administrative/metropolitan area. The innermost circle in dark green represents the core boundary (distance $r_{w}$ from the center $O_{w}$) and the two outer circles in lighter green denote the distance at which the occurrence probability $p_{w,c}$ is equal to one half and one third of the probability in the core $p_{0,w}$, respectively. The purple shaded area indicates the administrative area of each prefecture (top row) and the metropolitan employment area of each city (bottom row).
Motivated by these observations, our core-periphery model is formalized as follows. The occurrence probability of the binomial model is assumed to be constant within a certain distance from the center and decay as a power law outside of that range:
$p_{w,c}=\begin{cases}p_{0,w}&\text{for }d_{w,c}\leq r_{w},\\ p_{0,w}(d_{w,c}/r_{w})^{-a_{w}}&\text{for }d_{w,c}>r_{w}.\end{cases}$ (3)
This model has three free parameters: $p_{0,w}$ is the occurrence probability in cells within a radius of $r_{w}$ from the center, and $a_{w}$ denotes the exponent of the power-law decay outside the range.
We estimate the parameters of the space-dependent binomial model by maximizing the likelihood for the geographic distribution of each toponym. The fitted model shows, in general, a better agreement to the data in terms of the relationship between $n_{\mathrm{all}}$ and $n_{w}$, as shown in Fig 4B. For the results for all domestic toponyms studied in this work, see Fig S5 in Supporting Information. In particular, this model reproduces the two-branch scaling behavior of the empirical data. This significant improvement from the location-independent model is also quantitatively demonstrated by the decrease in Akaike information criterion (AIC) for all the toponyms studied in this work (Fig 7).
The advantage of the core-periphery model is that it conforms to an intuitive interpretation: the radius $r_{w}$ can be seen as the extent of the area to which the toponym refers, i.e. the core. Outside this core area, users refer to the place less often, but the decrease in probability is gradual as a function of distance and slow enough to be modeled as a power law (compared to, e.g., an exponential decay). To verify this interpretation, we compare the estimated core (the area within the estimated radius $\hat{r}_{w}$ from the center $O_{w}$) of each toponym with the extent of the administrative unit or metropolitan area to which it refers (Fig 6D). For all toponyms, the model identifies the center within the area to which each toponym refers, although the sizes of the cores vary and do not necessarily coincide with the administrative boundaries. For many toponyms, the core area detected by the model is smaller than the administrative area. This may indicate that users are more likely to make geotagged posts with these toponyms when they are in the central city of a prefecture or in the central area of a city. For “Fukuoka” and “Sendai”, the core has a geographic scale similar to the metropolitan area, suggesting that the use of these two toponyms is aligned with the extent of the corresponding metropolitan area. In the case of “Sapporo”, the core is larger than the metropolitan area. This could indicate that users tend to associate the toponym with a larger area; however, it is also possible that this is because the estimated parameter represents one of the many local maxima in the likelihood landscape.
[…]
Figure 7: Goodness of fit evaluated by Akaike information criterion (AIC).
IV Discussion and conclusion
In this article, we investigated the geographic patterns of toponym occurrence in social media using a dataset of geotagged Twitter posts. We found that the heterogeneous geographic distribution of geotagged posts is highly correlated with the population, especially with the employed population. This implies that the geotagged posts are, on the whole, representative of the language use in the population. The occurrence of each toponym in these geotagged posts is also characterized by geographic heterogeneity. Moreover, we found that the relationship between the total number of posts and the number of posts containing toponyms shows a distinctive scaling pattern. Comparison with patterns for common nouns and foreign toponyms suggests that this scaling pattern originates from the spatial specificity of toponym occurrence, which is successfully quantified by the dissimilarity from the location-independent model.
Finally, we presented the core-periphery model, which assumes a location-dependent occurrence probability of toponyms. Despite its simplicity with only three fitting parameters, this model can reasonably reproduce the empirically observed geographic distributions of toponym occurrence. This implies the following: First, each toponym has a core, i.e., a geographic area in which the toponym occurs with the highest probability, which can be regarded as the area that users collectively identify with the toponym. This interpretation is supported by the fact that the core is larger for the names of regions than for the names of cities and wards. Second, outside this core, the occurrence probability decreases slowly with distance following a power law. This may indicate that human attention, perception, and representation of geographic space respond nonlinearly to distance. It is also reminiscent of Tobler’s first law of geography: “everything is related to everything else, but near things are more related than distant things” [20]. In this context, our findings can be seen as another example of distance decay behavior in human activities, alongside other aspects such as commuting [21, 22, 23], tourism [24, 25, 26], and crime [27, 28, 29]. Yet, only few works have examined distance decay in linguistic patterns [30, 18], and we believe that our analysis is the first to confirm it on a large scale.
We note that our modeling approach does not aim to precisely replicate empirical observations of toponym occurrence. Rather, our models serve as a reference against which empirical data can be compared. As such, they simplify some aspects of the real-world toponym occurrence patterns. For example, in the core-periphery model, the model is isotropic, that is, the core has a circular boundary and the occurrence probability decreases uniformly in all directions. In reality, however, the area denoted by the toponym is not necessarily circular and may have elongated or irregular shapes, and the decay outside the core may not be isotropic due to geographical or transportation constraints. We also assumed no correlations between the occurrence of different toponyms, although there may be competitive or synergistic interactions between them that influence their occurrence patterns. It is because of these simplifications that our approach is able to provide an interpretable framework that allows us to identify the essential elements of the geospatial toponym distributions.
We also remark the possibility that geotagged posts on Twitter may not be an unbiased, representative sample of the language use of the general population. This issue can be divided into two questions: whether the geotagged Twitter users can be considered a good proxy for the population at large [31, 11, 32, 33], and whether the language use in geotagged Twitter posts is consistent with language use in other contexts [34, 35]. The effect of the first problem on our results is presumably relatively small compared to other work using geotagged Twitter posts to study language use, for two reasons: (1) we focus only on the content of the posts without correlating them with user demographics, and (2) the age and socioeconomic status of users are unlikely to significantly affect their use of toponyms. It is however possible that urban toponyms are overrepresented due to population biases, which could affect the geographic distribution of toponyms. In this work, we focused on relatively large cities and wards within them, but whether our findings generalize to the names of smaller towns and villages needs to be carefully examined in future research. The second question concerns the generalizability of our findings to the use of toponyms in other contexts. To answer this question, one needs to analyze data from other social media platforms or in other forms of communication. Although this is an important point, it is beyond the scope of our study and should be addressed in future work.
Acknowledgements.
T.H. acknowledges the computational resources provided by the Aalto Science-IT project. N.F. was supported by JSPS KAKENHI Grant Number 24K03007 and JST PRESTO Grant Number JPMJPR21RA, Japan. Data availability: All data and code necessary to reproduce the findings of this paper are deposited and publicly available at https://doi.org/10.5281/zenodo.13860969 and https://github.com/takayukihir/geotagged-tweets. Due to the Twitter’s terms of service, we are unable to redistribute the raw text of the posts we have obtained from the Twitter API. Instead, the data on the number of geotagged posts aggregated at the level of basic grid square cells are available in the above repository. The resident and employed population data are based on 2015 Population Census and 2016 Economic Census for Business Activity, respectively, and are available from the Statistics Bureau of Japan (https://www.stat.go.jp/english/data/kokusei/2015/summary.html, https://www.stat.go.jp/english/data/e-census/2016/outline.html).
Supporting information
S1 Definition of relative entropy
In the main text, we use relative entropy (Kullback–Leibler (KL) divergence) to measure the dissimilarity between the empirical and model distributions. Here, we describe the definition of relative entropy in detail.
Each distribution is defined in the two dimensional space of the total number of posts per cell, denoted by $n_{\mathrm{all}}$, and the number of posts with word $w$, denoted by $n_{w}$. Specifically, we only consider the values of $n_{\mathrm{all}}$ that are empirically observed, and the values of $n_{w}$ equal to or smaller than $n_{\mathrm{all}}$. That is, the support of each distribution is given by
$\{(n_{\mathrm{all}},n_{w})\in\mathbb{N}\times\mathbb{N}\mid(\exists c\in\Gamma)[n_{\mathrm{all}}=n_{\mathrm{all},c}],n_{w}\leq n_{\mathrm{all}}\}$
for a set of grid cells $\Gamma$. Given a set of empirical numbers of posts $\{n_{w,c}\}$ containing word $w$ for $c\in\Gamma$, relative entropy is defined as
$D_{\mathrm{KL}}(Q_{w}\parallel P_{w})=\frac{1}{|\Gamma|}\sum_{c\in\Gamma}\left[\log Q_{w}(n_{w,c}\mid n_{\mathrm{all},c})-\log P(n_{w,c}\mid n_{\mathrm{all},c},\,\hat{p}_{w})\right],$ (S1)
where $P_{w}:=P(n_{w}\mid n_{\mathrm{all}},\hat{p}_{w})$, and $Q_{w}(n_{w}\mid n_{\mathrm{all}})$ denotes the empirical distribution of $n_{w}$ for given $n_{\mathrm{all}}$ over the set of grid cells $\Gamma$.
In Fig S1A, we present the estimated relative entropy for each toponym and each noun. As expected, toponyms denoting cities (such as “Nagoya”) and some submunicipal districts (such as “Shinjuku”), as well as place nouns (such as “park”), are characterized by relatively large values of relative entropy than those for common nouns without a place connotations (e.g., “wallet”), implying that the empirical distribution is highly dissimilar from the location-independent binomial model. However, the relative entropies for toponyms that represent regions (e.g., “Shikoku”), prefectures (e.g., “Ishikawa”), and some other submunicipal districts (e.g., “Setagaya”) are as small as those for common nouns. those for common nouns without a place connotations (e.g., “wallet”). This seems inconsistent with what we expect from S3 Fig in Supporting Information, in which the difference between the data and the model is clearly visible for all domestic toponyms. This is presumably because the data contain more grid cells without any toponym occurrence, i.e., $n_{w}=0$, than expected by the model. Such cells are prevalent even in cells with large $n_{\mathrm{all}}$, in which the model predicts a low probability for $n_{w}=0$. The increase in relative entropy due to the presence of such zero-occurrence grid cells will be more pronounced for words with a large $\hat{p}_{w}$, or, equivalently, a large number of posts $N_{w}$. This hypothesis is supported by the strong positive correlation between $N_{w}$ and $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$, as shown in Fig S1B. Simply put, a large value of relative entropy may just be an artifact of the popularity of the word.
Figure S1: Dissimilarity of empirical data from the location-independent model, evaluated by relative entropy. (A) Relative entropy $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$ for each toponym $w$ and noun studied in this work. For each word, 20000 grid cells are randomly sampled 50 times to compute the mean and 95% confidence intervals, indicated by error bars. (B, C) Relative entropy $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$ and the modified version $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$, plotted against the total number of posts $N_{w}$ containing word $w$.
To discount the effect of inflated zero occurrences, we only sample grid cells with non-zero occurrences for computing the empirical distribution and relative entropy, i.e., $\Gamma\subseteq\{c\mid n_{w,c}>0\}$. Let $\tilde{Q}$ denote the empirical distribution thus obtained. As shown in Fig 5 in the main text, this approach clearly discriminates the domestic toponyms and place nouns from nonspatial nouns and foreign toponyms. The correlation between $N_{w}$ and $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$ is much weaker (Fig S1C), further corroborating the hypothesis that the value of $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$ is dominated by the contribution from zero-occurrence grid cells.
Figure S2: Probability density functions of occurrence density $\sigma_{w}$.
Figure S3: Scatter plot of $n_{w}$ versus $n_{\mathrm{all}}$ for all toponyms and nouns studied in this work.
Figure S4: Comparison between the empirical data and the location-independent model. Each set of contours represents the kernel density plot of $n_{w}$ versus $n_{\mathrm{all}}$ of empirical data (blue) and the location-independent model (red).
Figure S5: Comparison between the empirical data and the core-periphery model. Each set of contours represents the kernel density plot of $n_{w}$ versus $n_{\mathrm{all}}$ of empirical data (blue) and the core-periphery model (red).